# Grounded Test-Time Adaptation for LLM Agents

### v1

@@ 1 Introduction
Large language model (LLM)-based AI agents have demonstrated remarkable capabilities across diverse tasks (li_can_2023; zhou_webarena_2024; patil_berkeley_nodate; xie_osworld_2024), yet they often fail to generalize when deployed in novel agentic environments such as web navigation (zhou_webarena_2024; he_webvoyager_2024) and function calling (patil_berkeley_nodate; yao_tau-bench_2024). These failures reflect a core challenge—the systematic mismatch between the agent’s pre-trained knowledge and the specific syntax and dynamics of the deployed environment (yang_agentoccam_2025; thil_navigating_2024). We classify the mismatch manifested in two primary failure modes, which we illustrate by considering an LLM agent tasked with booking a flight on a previously unseen travel website:
@@
To address the the two failure modes and close the generalization gap, we propose two annotation-free strategies that operate at deployment time. (1) A parametric online adapter (parametric test-time adaptation) — a lightweight module that learns a small per-interaction bias (an adaptation vector applied to late features) to quickly align an agent’s output distribution with environment-specific syntax. Observing Go and dest_field on the page, for example, lets the adapter steer action generation toward the site’s actual element names. (2) A non-parametric dynamics grounding (non-parametric test-time adaptation) — a short, persona-guided exploration phase that iteratively collects in-context rules about state transitions (e.g., clicking Go opens a date pop-up) and supplies those rules as context for downstream planning. Both methods require only test-time observations and no trajectory annotations or expensive fine-tuning.
@@ 2 Related Work
Originating in computer vision, test-time adaptation (TTA) addresses distributional shifts between training and testing data by adapting models at inference time (wang_tent_2021; niu_efficient_2022; sun_test-time_2020). Our parametric test-time adaptation builds directly on methods that update model parameters or steering vectors at test time to minimize an unsupervised objective, such as entropy minimization or self-supervised learning (wang_tent_2021; hu_slot_2025; li_inference-time_2023). While TTA for LLMs is emerging for tasks like math reasoning (zuo_ttrl_2025) and few-shot learning (akyurek_surprising_2025), its application to the unique, interactive, and stateful challenges of LLM agents remains underexplored. Our work is the first to systematically apply this paradigm to adapt agents to novel environment formats without supervised trajectories.
@@ 3.1 Parametric Test-Time Adaptation
Figure 1: Overview of parametric test-time adaptation. (1) At the start of each episode, we initialize an adaptation vector $\delta$ as a zero vector. (2) During task execution, the agent receives environment observations and takes actions, with the adaptation vector applied as a bias to the final hidden layer. (3) At each step, we update the adaptation vector using cross-entropy loss on the current input, enabling rapid alignment to environment-specific observation and action formats. (4) The updated vector shifts the model’s output distribution to better match the test-time environment, correcting syntactic mismatches without full fine-tuning.
To address the syntactic mismatches that arise in novel environments, we propose parametric test-time adaptation. The objective is to parametrically adapt the model’s output distribution at test time by treating the current context (i.e. task instructions, observations, and action history) as a self-supervisory signal. This allows the model to align with local syntactic patterns, such as unique UI element labels or response formats, without requiring a deep understanding of the environment’s causal dynamics.
@@
This vector $\delta$ is updated at each turn by performing one step of gradient descent to minimize the standard cross-entropy language modeling loss on the current input context, $\mathcal{I}$:
@@
where $f_{\theta,\delta}$ is the model parameterized by its fixed weights $\theta$ and the adaptable vector $\delta$, and $\eta$ is the learning rate. This update encourages the model to internalize the syntactic features present in the immediate context.
To see this in practice, consider our agent on the unseen website travel-example.com. The agent’s input context $\mathcal{I}$ contains the accessibility tree, which includes strings such as [145] <button>Go and dest_field.
@@
This process is computationally efficient and step 2-4 repeats at every step within an episode. To prevent catastrophic forgetting and ensure adaptation is specific to the current environment, $\delta$ is reset to zero at the beginning of each new episode. This lightweight update aligns the model’s pre-trained distribution with the specific distribution of the test-time environment.
@@ 3.2 Non-parametric Test-Time Adaptation
Figure 2: Overview of non-parametric test-time adaptation. (1) We synthesize diverse exploration tasks based on personas using environment descriptions. (2) An exploration agent interacts with the environment to collect interaction logs of state transitions. (3) An LLM extracts and summarizes environment dynamics from these logs. (4) A reasoning model filters these dynamics, which are then used to augment the agent’s context during evaluation, enabling more transition-aware decision making.
LLM-based agents often fail in novel environments due to a semantic mismatch: they lack a “world model” to predict the outcomes of their actions (ha_world_2018; chae_web_2025). To address this, we propose non-parametric test-time adaptation, a strategy that proactively discovers an environment’s causal dynamics before task execution. The objective is to construct a non-parametric, in-context world model by performing a one-time, automated exploration. The cost of this discovery phase is amortized over all subsequent tasks performed in that environment.
@@ 4.1 Setup
Table 2: Main results on WebArena, WebVoyager, BFCLv3, and Tau-Bench benchmarks. We report task success rates (%) for each model and adaptation method. For Tau-Bench, we average runs across 5 seeds and use a custom more stable codebase. Both parametric (PA) and non-parametric (NPA) test-time adaptation improve performance.
@@
Qwen2.5-14B-Instruct (+ NPA) 20.0 (+3.0) 22.0 (+3.5) - -
Qwen2.5-14B-Instruct (hybrid) 21.0 (+4.0) 21.0 (+2.5) - -
@@ 4.2 Results and Analysis
Table 4: Relative latency gain (%) of parametric test-time adaptation with different number of training steps on BFCLv3 multi-turn compared to the baseline (no adaptation). Learning rate is fixed to $0.1$.
@@ 5 Limitations and Future Work
Our study is primarily focused on the Qwen2.5 model family for parametric adaptation; future work should validate these findings across a broader range of open-source architectures. Additionally, our online parametric adaptation method does not currently normalize updates by the hidden dimension size, which could improve robustness to hyperparameters.

### v2

@@ 1 Introduction
Large language model (LLM)-based AI agents have demonstrated remarkable capabilities across diverse tasks (li_can_2023; zhou_webarena_2024; patil_berkeley_nodate; xie_osworld_2024), yet they often fail to generalize when deployed in novel agentic environments such as web navigation (zhou_webarena_2024; he_webvoyager_2024) and function calling (patil_berkeley_nodate; yao_tau-bench_2024). These failures reflect two core challenges—the systematic mismatch between the agent’s pre-trained knowledge and the specific syntax and dynamics of the deployed environment (yang_agentoccam_2025; thil_navigating_2024). While both mismathes stem from a lack of prior knowledge, we distinguish them operationally because they require different adaptation mechanisms (parametric vs. non-parametric). To illustrate these two failure patterns, we consider an LLM agent tasked with booking a flight on a previously unseen travel website:
@@
To address the the two failure modes and close the generalization gap, we propose two annotation-free strategies that operate at deployment time. (1) A parametric online adapter (parametric test-time adaptation)–a lightweight module that learns a small per-interaction bias (an adaptation vector applied to late features) to quickly align an agent’s output distribution with environment-specific syntax. Observing Go and dest_field on the page, for example, lets the adapter steer action generation toward the site’s actual element names. (2) A non-parametric dynamics grounding (non-parametric test-time adaptation)—a short, persona-guided exploration phase that iteratively collects in-context rules about state transitions (e.g., clicking Go opens a date pop-up) and supplies those rules as context for downstream planning. Both methods require only test-time observations and no trajectory annotations or expensive fine-tuning.
@@ 2 Related Work
Originating in computer vision, test-time adaptation (TTA) addresses distributional shifts between training and testing data by adapting models at inference time (wang_tent_2021; niu_efficient_2022; sun_test-time_2020). Our parametric test-time adaptation builds directly on methods that update model parameters or steering vectors at test time to minimize an unsupervised objective, such as entropy minimization or self-supervised learning (wang_tent_2021; hu_slot_2025; li_inference-time_2023). Our parametric test-time adaptation aligns with steering vector research (tan_analyzing_2025; sinii_steering_2025; panickssery_steering_2024; turner_steering_2024) in its use of latent space biases to shift output distributions. However, standard steering approaches generally apply modulation vectors to shift high-level traits  (tan_analyzing_2025), such as honesty (zou_representation_2025). We distinguish our work by tailoring this mechanism to the dynamic nature of agentic tasks: our vector is treated as a temporary parameter, updated online using the environment’s response as a self-supervisory signal, and reset per episode to ensure precise, context-dependent alignment without interference across tasks. While TTA for LLMs is emerging for tasks like math reasoning (zuo_ttrl_2025) and few-shot learning (akyurek_surprising_2025), its application to the unique, interactive, and stateful challenges of LLM agents remains underexplored. Our work is the first to systematically apply this paradigm to adapt agents to novel environment formats without supervised trajectories.
@@ 3.1 Parametric Test-Time Adaptation
Figure 1: Overview of parametric test-time adaptation. This figure includes an example of web navigation shopping task to illustrate how the agent adapts to new environment. (1) At the start of each episode, we initialize an adaptation vector $\delta$ as a zero vector and construct inputs to the LLM agent. (2) During task execution, the agent receives environment instructions and observations. (3) At each step, we update the adaptation vector using cross-entropy loss on the current input, and apply the adaptation vector as a bias to the LLM’s final hidden layer. This enables rapid alignment to environment-specific observation and action formats. (4) The LLM agent takes a new action with the updated vector, which shifts the model’s output distribution to better match the test-time environment.
To address syntactic mismatches that arise in novel environments, we propose parametric test-time adaptation. The objective is to parametrically adapt the model’s output distribution at test time by treating the current context (i.e. task instructions, observations, and action history) as a self-supervisory signal. This allows the model to align with local syntactic patterns, such as unique UI element labels or response formats, without requiring a deep understanding of the environment’s causal dynamics.
@@
This vector $\delta$ is updated at each turn by performing one gradient descent step to minimize the language modeling loss of the current input context $\mathcal{I}$ according to the following:
@@
where $f_{\theta,\delta}$ is the LLM parameterized by its fixed weights $\theta$ and the adaptable vector $\delta$, and $\eta$ is the learning rate. $\mathcal{I}_{1:n-1}$ denotes the input subsequence consisting of tokens 1 through n-1, and $\mathcal{I}_{2:n}$ denotes the corresponding next-token targets obtained by shifting the sequence by one position. The loss in Eq. 3 is therefore computed by predicting each token in $\mathcal{I}_{2:n}$ from its preceding context $\mathcal{I}_{1:n-1}$. This update encourages the model to internalize the syntactic features present in the immediate context by adapting its output distribution to the environment context.
To illustrate this in practice, consider our agent on the unseen website travel-example.com. The agent’s input context $\mathcal{I}$ contains the accessibility tree, which includes strings such as [145] <button>Go and dest_field.
@@
This process is computationally efficient, and step 2-4 is repeated at every model prediction within an episode. To prevent catastrophic forgetting and ensure that adaptation is specific to the current environment, $\delta$ is reset to zero at the beginning of each new episode. This lightweight update aligns the model’s pre-trained distribution with the specific distribution of the test-time environment.
@@ 3.2 Non-parametric Test-Time Adaptation
Figure 2: Overview of non-parametric test-time adaptation. This figure includes an example of web navigation shopping task to illustrate how the pipeline generates environment dynamics in language. (1) We synthesize diverse exploration tasks based on personas using environment descriptions. (2) An exploration agent interacts with the environment to collect interaction logs of state transitions and updates the environment rules/dynamics dynamically. (3) An LLM extracts and summarizes environment dynamics from these logs. (4) A reasoning model filters less informative rules, which are then used to augment the agent’s context during evaluation, enabling more transition-aware decision making.
LLM-based agents often fail in novel environments due to a semantic mismatch: they lack a “world model” to predict the outcomes of their actions (ha_world_2018; chae_web_2025). To address this, we propose non-parametric test-time adaptation, a strategy that proactively discovers the causal dynamics of an environment before the execution of the task. The objective is to construct a non-parametric (in language), in-context world model by performing a one-time, automated exploration. The cost of this discovery phase is amortized over all subsequent tasks performed in that environment.
@@ 4.1 Setup
Table 2: Main results on WebArena, WebVoyager, BFCLv3, and Tau-Bench benchmarks. We report task success rates (%) for each model and adaptation method. For Tau-Bench, we average runs across 5 seeds and use a custom more stable codebase. Both parametric (PA) and non-parametric (NPA) test-time adaptation improve performance. ”N/A” indicates not applicable—AWM requires training a web-specific world model on state transitions, which is not applicable to Tau-Bench (conversational only) or BFCLv3 (no explicit web-like state transition data); Non-parametric adaptation does not operate in conversational Tau-Bench as there is no explicit and fixed state transition rules in the environment.
@@
Qwen2.5-14B-Instruct (+ NPA) 20.0 (+3.0) 22.0 (+3.5) N/A N/A
Qwen2.5-14B-Instruct (hybrid) 21.0 (+4.0) 21.0 (+2.5) N/A N/A
@@ 4.2 Results and Analysis
Table 4: Relative latency gain (%) of PA on BFCLv3 multi-turn. LR fixed to $0.1$.
@@ 5 Limitations and Future Work
Our study is primarily focused on the Qwen2.5 model family for parametric adaptation; future work should validate these findings across a broader range of open-source architectures. Our non-parametric adaptation method currently operates in environment with explicit state transitions, where testbeds such as user-oriented conversational task is not explored. Additionally, our online parametric adaptation method does not currently normalize updates by the hidden dimension size, which could improve robustness to hyperparameters.